# An LLM-based Agent for Reliable Docker Environment Configuration

### v2

@@ A.4 Repository assignment
A.5 Results
Based on the times shown in the logs, we calculated that the average configuration time for each repository is 21.33 minutes. Furthermore, none of the repositories that Repo2Run failed to configure were successfully configured manually. Additionally, in the manual experiment, five environments that were successfully configured by Repo2Run were not successfully configured, representing 15.6% of the total successfully configured assignments.
Appendix B Safe commands
Appendix C Repo2Run tools
@@
Table 6: Command list and their functions
@@
waitinglist add -p package_name Add item into waiting list. If no “version_constraints” are
[-v version_constraints] -t tool specified, the latest version will be downloaded by default.
waitinglist addfile file_path Add all entries from a file similar to requirements.txt format to
the waiting list. Format should be package_name [version_constraints].
@@
conflictlist solve -v Resolve the conflict for the first element in the conflict list, and
[version_constraints] update the version constraints for the corresponding package_name and tool to version_constraints. If no “version_constraints” are specified, the latest version will be downloaded by default. The package_name and tool in the original waiting list must match one of the elements in the conflictlist. Here, the version_constraints are specified.
@@
runtest Check if the configured environment is correct using“pytest”.
poetryruntest Check if the configured environment is correct in the poetry environment. If you want to run tests in the poetry environment, run it.
@@
change_python_version Switching the Python version in the Docker container will forgo
python_version any installations made prior to the switch. The Python version number should be represented directly with numbers and dots, without any quotation marks.
clear_configuration Reset all the configuration to the initial setting of python:3.10.
Appendix D Benchmark
@@
Appendix E pipreqs baseline settings

### v3

@@ E.1.4 Repository assignment
E.1.5 Results
Based on the times shown in the logs, we calculated that the average building time for each repository is 21.33 minutes. Furthermore, none of the repositories that Repo2Run failed to build were successfully built manually. Additionally, in the manual experiment, five environments that were successfully built by Repo2Run were not successfully built, representing 15.6% of the total successfully built assignments.
The table information and logs are too lengthy, so we place them on https://anonymous.4open.science/r/Repo2Run.
E.2 Industry-level repository building validation
As shown in Table LABEL:tab:industry_evaluation, we present the evaluation of Repo2Run  on 59 popular industry-grade repositories. Additionally, Table LABEL:tab:industry_evaluation1 shows the successful configuration of additional industry-grade repositories.
Table 11: Evaluation of Repo2Run on 59 popular industry-grade repositories. The repositories were selected from GitHub Ranking [48], filtering out those without tests and with #LOC < 10,000.
Configuration state #LOC > 10,000 #LOC > 100,000 Medium Max
Success(30) 30 (100%) 14 (46.7%) 96,157 2,389,065
Fail(29) 29 (100%) 24 (82.6%) 333,464 2,669,973
All(59) 59 (100%) 40 (9.3%) 124,352 2,669,973
Table 12: Successful additional configuration of 30 industry-grade repositories.
Repository #LOC Repository #LOC Repository #LOC
comfyanonymous/ ComfyUI 737,543 OpenBB-finance/OpenBB 2,389,065 scrapy/scrapy 77,278
yt-dlp/yt-dlp 155,662 keras-team/keras 204,537 AUTOMATIC1111/ stable-diffusion-webui 40,448
pallets/flask 25,190 geekan/MetaGPT 82,305 psf/black 118,551
deepfakes/faceswap 74,690 streamlit/streamlit 201,894 psf/requests 12,782
lllyasviel/Fooocus 328,199 labmlai/annotated_ deep_learning _paper_implementations 510,851 RVC-Boss/GPT-SoVITS 34,658
mingrammer/ diagrams 11,994 OpenInterpreter/open-interpreter 15,818 hiyouga/LLaMA-Factory 76,000
nvbn/thefuck 10,267 pandas-dev/pandas 571,849 ytdl-org/youtube-dl 96,157
TheAlgorithms/ Python 419,917 fastapi/fastapi 177,120 Textualize/rich 45,744
All-Hands-AI/OpenHands 160,514 Stability-AI/ stablediffusion 13,698 ultralytics/ultralytics 94,201
gradio-app/gradio 184,722 QuivrHQ/quivr 10,264 freqtrade/freqtrade 188,889
Appendix F Repo2Run tools
@@
Table 13: Command list and their functions.
@@
waitinglist add -p package_name [-v version_constraints] -t tool Add item into waiting list. If no “version_constraints” are specified, the latest version will be downloaded by default.
waitinglist addfile file_path Add all entries from a file similar to requirements.txt format to the waiting list. Format should be package_name [version_constraints].
@@
conflictlist solve -v[version_constraints] Resolve the conflict for the first element in the conflict list, and update the version constraints for the corresponding package_name and tool to version_constraints. If no “version_constraints” are specified, the latest version will be downloaded by default. The package_name and tool in the original waiting list must match one of the elements in the conflictlist. Here, the version_constraints are specified.
@@
runtest Check if the built environment is correct using“pytest”.
poetryruntest Check if the built environment is correct in the poetry environment. If you want to run tests in the poetry environment, run it.
@@
change_python_version python_version Switching the Python version in the Docker container will forgo any installations made prior to the switch. The Python version number should be represented directly with numbers and dots, without any quotation marks.
clear_configuration Reset all the buildings to the initial setting of python:3.10.
Appendix G Repository building success status
As shown in Table LABEL:tabs:success_status, we list the success status of all the packages in our constructed benchmark.
@@ Appendix G Repository building success status
Appendix H Baseline settings
H.1 Pipreqs settings
@@
H.2 SWE-agent settings
Below is the settings of SWE-agent for building the environment:
agent:
templates:
system_template: |-
SETTING: You are an expert skilled in environment configuration, operating directly in the command line with a specialized interface.
The specialized interface consists of {{WINDOW}} lines of context at a time. In addition to typical bash commands, you can also use the following commands to help you with environment setup.
COMMANDS:
{{command_docs}}
Please note that THE EDIT COMMAND REQUIRES PROPER INDENTATION. Ensure correct indentation when modifying code.
RESPONSE FORMAT:
Your shell prompt is formatted as follows:
(Current task: <task name>) <cwd> \$
You need to format your output using two fields: discussion and command.
Example:
DISCUSSION
Ill start by using ls to see the files in the root directory.
ls /repo
You should only include a SINGLE command in the command section and then wait for a response. Everything in the DISCUSSION section will be saved for future reference.
You can use any other bash commands and the special commands listed above, but interactive session commands (e.g., python, vim) are not supported.
instance_template: |-
Were currently setting up the environment for the following task. Here are the details:
TASK:
{{task_details}}
INSTRUCTIONS:
Now, youll carry out this task on your own. Your terminal session has begun in the repositorys root directory. Use the provided commands and any bash commands you need. Edit and check files as needed.
The goal is to generate a Dockerfile that can successfully build and run the tests in the repository using the command "pytest --collect-only -q".
NOTE:
1. The repository is cloned into /repo.
2. The Dockerfile should start with the following lines (if the base image is python:3.10 and repository is adamobeng/wddbfs):
FROM python:3.10
RUN pip install pytest
RUN git clone https://github.com/adamobeng/wddbfs.git
RUN mkdir /repo
RUN git config --global --add safe.directory /repo
RUN cp -r /wddbfs/. /repo && rm -rf /wddbfs/
RUN rm -rf /wddbfs
Your task includes:
0. **Generate Dockerfile**: Create a file named "Dockerfile" in the root path (e.g, the absolute path is /Dockerfile).
1. **Read Directory Structure**: Check the folder structure in the root directory.
2. **Check the Configuration Files**: Inspect files like "requirements.txt", "setup.py", "setup.cfg", "Pipfile*", etc.
3. **Determine Package Dependencies**: Handle dependencies and manage conflicting dependency versions.
4. **Testing**: Ensure "pytest /repo --collect-only -q" runs without errors.
5. **Generate Dockerfile**: Write necessary installation or setup steps determined from the inspection. If you finish run testing successfully, you should modify the Dockerfile in /Dockerfile.
IMPORTANT TIPS:
* Check the directory and files carefully.
* Make sure Dockerfile commands are correct.
* Use proper Dockerfile syntax and indentation.
* Test the final Dockerfile by running "pytest /repo --collect-only -q".
(Current task: Generate Dockerfile)
(Current directory: {{working_dir}})
bash-\$
next_step_template: |-
{{observation}}
(Current task: Generate Dockerfile)
(Current directory: {{working_dir}})
bash-\$
next_step_no_output_template: |-
Your command ran successfully and did not produce any output.
(Current task: Generate Dockerfile)
(Current directory: {{working_dir}})
bash-\$
tools:
env_variables:
WINDOW: 100
OVERLAP: 2
bundles:
- path: tools/registry
- path: tools/defaults
- path: tools/search
- path: tools/edit_linting
- path: tools/submit
- path: tools/env_setup
parse_function:
type: thought_action
history_processors:
- type: last_n_observations
n: 5
Appendix I Failure Case
Figure 12 illustrates a “module not found” error when building the repository “jialuechen/deepfolio”. In this case, the issue arises from the absence of updating the unit tests. The test file “test_stats.py” attempts to import modules from “deepfolio.stats”, but “stats” does not exist in “deepfolio”. Consequently, no matter how the LLM agent operates, it cannot directly run this test. This highlights the importance for developers to continuously update existing unit tests as the code repository evolves.